Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Tue Feb 24 09:47:47 MET 1998) Minor failures
The first 1.5 hours (till 8:12) lost due to wrong observing frequency (C-band) set. Then, the last 2 minutes of the scan at 11:49 not recorded because (fortunately) the low tape sensor switched off the tape drive preventing (scheduled!) complete unrolling of the tape. The sensivity of the system was poorer than usually; reason as yet unknown.

Westerbork (tony foley at Tue Feb 24 14:51:51 MET 1998) Success
Observation looks OK

Noto (C.Trigilio at Wed Feb 25 14:35:20 MET 1998) Success

Jodrell (Alastair Gunn at Thu Feb 26 14:52:26 MET 1998) Success
Strong RFI, receiver often saturated. No other known problems.

Urumqi (Liu Xiang at Fri Feb 27 05:02:35 MET 1998) Success

Medicina (M.Roma at Fri Feb 27 10:40:28 MET 1998) Success
No problems

Onsala (A. Polatidis at Mon Mar 2 14:43:04 MET 1998) Minor failures
Well major failures. There has been a leak in the receiver causing Tsys to be ~70% higher. Telescope was stowed between 0857 and 1432 due to high winds. At 1458 the inchworm motor for head2 failed and it took us until the end of the experiment to fix the INCHWORM CONTROL MODULE. Sorry about this.

Calibration (jean\-francois Desmurs at Tue Mar 3 11:56:34 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name em25b.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 16:09:04 MET 1998) Minor failures
missed 0630-0636 and 0916-0922. Around 1321 antenna trouble.